The top of a rectangular dining room table is 4 times as long as it wide. Find the dimensions of the table if the area is 36 ft.
Answer by lwsshak3(11628)

The top of a rectangular dining room table is 4 times as long as it wide. Find the dimensions of the table if the area is 36 ft.
***
let x=width of table
4x=length of table
width*length=area
x*4x=36
4x^2=36
x^2=9
x=√9=3
4x=12
width of table=3 ft
length of table=12 ft